At what temperature will the root-mean-square speed of oxygen molecules have the value of 640 m/s? 1 kilomole of oxygen has a mass of 32kg.
1
Expert's answer
2013-01-14T10:52:34-0500
v =sqrt(3RT/m)
where m is molar mass (in out case 0.032 kg)
R is gas constant = 8.3
hence we have
T = v^2*m/(3R) =640^2*0.032/(3*8.3) = 526.4 K = 253.25 C